OpenStudy (anonymous):

Four more than twice a number is seven more than the number. What is the number?

OpenStudy (anonymous):

4 more than twice a number is seven more than the number 3x2=6+4=10 so the number is 3

OpenStudy (anonymous):

7+3=10

OpenStudy (texaschic101):

4 more then twice a number .... 2n + 4 is means = 7 more then the number...n + 7 2n + 4 = n + 7

OpenStudy (anonymous):

A. 2003-05-03-00-00_files/i0160000.jpg B. 1 C. 3 D. 2003-05-03-00-00_files/i0160001.jpg